Mr and Mrs. Seller want to net a 12% profit after paying the brokerage firm 6.5% commision. If the original purchase price was $104,500, what is the miniumum offer they can accept?
Answer #1

The seller wants a profit of 12%, so the total amount including only 12% profit=104500+12%*104500=117040
Now, commission is 6.5%
=>100%-6.5%=93.5%

Minimum offer to accept=117040/93.5%
=$125176.4706 or $125176 (Rounded to the nearest dollar)